OpenStudy (anonymous):

How do you write the number 5.869 × 1011 in standard form?

OpenStudy (solomonzelman):

go ahead.

OpenStudy (solomonzelman):

go ahead, ninja

OpenStudy (solomonzelman):

oh she left

OpenStudy (solomonzelman):

you mean \(\large\color{slate}{ 5.869\times 10^{11}}\) ?

OpenStudy (anonymous):

yeah sorry lol

OpenStudy (solomonzelman):

for any integer c, where \(\large\color{slate}{ c\ne0}\). \(\large\color{teal}{ \text{_}{\scriptsize~}.\text{_}{\scriptsize~}\text{_}{\scriptsize~}\text{_}\times 10^{\LARGE c}}\) move the decimal c spouts to the right. \(\large\color{teal}{ \text{_}{\scriptsize~}.\text{_}{\scriptsize~}\text{_}{\scriptsize~}\text{_}\times 10^{\LARGE -c}}\) move the decimal c spouts to the left.

OpenStudy (anonymous):

i have no idea what that means lol

OpenStudy (anonymous):

i kinda suck at math

OpenStudy (solomonzelman):

I will give you examples

OpenStudy (solomonzelman):

I also suck at math

OpenStudy (solomonzelman):

For example: \(\large\color{teal}{ 12{\scriptsize~}.3\times 10^{\LARGE -2}}\) >>> \(\large\color{teal}{ 0.{\scriptsize~}123}\) \(\large\color{teal}{ 56{\scriptsize~}.41234\times 10^{\LARGE 4}}\) >>> \(\large\color{teal}{ 564123.{\scriptsize~}4}\)

OpenStudy (anonymous):

ok i get it thanks

OpenStudy (solomonzelman):

yw:)

OpenStudy (solomonzelman):

you are just moving the decimal.

OpenStudy (solomonzelman):

can you show me what you get, to verify your answer/